3. Check which of the following are solutions of the equation x - 2y = 4 and which are not:

(i) (0, 2) (ii) (2, 0) (iii) (4, 0)
(iv) ( ../2 , 4 ../2) (v) (1, 1).

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Consider the equation x - 2y = 4 ...(A) (i) For (0, 2), substituting x = 0, y = 2 in (A), we get 0 - 4 = 4 => - 4 = 4, not true. Hence, (0, 2) is not a solution. (ii) For (2, 0), substituting x = 2, y = 0 in (A), we get 2 - 0 = 4 => 2 = 4, not true. Hence, (2, 0) is not a solution. (iii) For (4, 0), substituting x = 4, y = 0 in (A), we get 4 - 0 = 4 => 4 = 4, true. Hence, (4, 0) is a solution. (iv) For ( J2 , 4../2 ), substituting x = .J2 , y = 4 .J2 in (A), we get • .J2 - 8.J2 = 4 => - 7 J2 = 4, not true. Hence, ( J2 , 4 ../2 ) is not a 8olution. (v) For (1, 1), substituting x = 1, y = 1 in (A), we get • 100 1- 2 = 4 => - 1= 4, not true. Hence, (1, 1) is not a solution.